### v1

# Expansion dynamics of strongly correlated lattice bosons

Abstract
We study the spatio-temporal dynamics of interacting bosons on a two-dimensional Hubbard lattice in the strongly interacting regime, taking into account the dynamics of condensate amplitude as well as the direct transport of non-condensed fluctuations. To that end we develop a selfconsistent density-matrix approach which goes beyond the standard Gutzwiller mean-field theory. Starting from the Liouville-von-Neumann equation we derive a quantum master equation for the time evolution of the system’s local density matrix at each lattice site, with a dynamical bath that represents the rest of the system. We apply this method to the expansion dynamics of an initially prepared cloud of interacting bosons in an optical lattice. We observe a ballistic expansion of the condensate, as expected, followed by slow, diffusive transport of the normal bosons. We discuss, in particular, the robustness of the Mott insulator phase as well as its melting due to incoherent transport. The method should be applicable to various models of lattice bosons in the strongly correlated regime.
I Introduction
The experimental realization of the superfluid-to-Mott insulator (SF-MI) transition in the Bose-Hubbard model (BHM) opened up new avenues to study non-equilibrium dynamics of strongly interacting Bose gases [1]. Since then, a variety of correlated models have been realized in the cold atom platform [2]. Moreover, the invention of quantum-gas microscopes has enabled the investigation of the spatio-temporal evolution of lattice bosons and their correlation properties in different dimensions as well as for different interaction strengths [3, 4]. The experimental observation of bimodal expansion of interacting Bose gases with ballistic and diffusive components [5, 6, 5] calls for a detailed understanding of these phenomena, involving the spatio-temporal interplay between the phase-coherent condensate dynamics and the incoherent excitations, especially in the strongly correlated regime.
The quench dynamics of Bose-Hubbard systems in one dimension (1D) are well described by, for example, the time-dependent density matrix renormalization group (t-DMRG) and its recent variant, the matrix product states approach [7, 8, 9]. Time-dependent exact diagonalization (ED) [8] and variational quantum Monte Carlo (t-VMC) [10] techniques are limited by short evolution times and finite system size and, thus, the difficulty to describe spontaneous symmetry breaking. However, due to the absence of long-range order in 1D a true condensation cannot occur [11]. While true symmetry-breaking with the formation of a condensate can occur in two and higher dimensional systems, describing the spatio-temporal dynamics of those systems has remained a challenge. Selfconsistent, perturbative quantum field-theoretical methods were developed for the homogeneous temporal dynamics [12] and to capture the coupled dynamics of condensate and non-condensed fluctuations in the weakly interacting regime [13, 14], which does not capture the Mott phase. The strongly interacting regime, on the other hand, is commonly described by time-dependent Gutzwiller methods based on single-site [15, 16] or cluster mean-field theory [17, 18]. This type of mean-field theory, however, neglects quantum hopping to neighboring sites and, thus, cannot describe temporal expansion initiating from a Mott localized state. In bosonic dynamical mean-field theory (B-DMFT), such hopping processes are included at second order [19]. It has been applied to describe the non-equilibrium dynamics of Bose-Hubbard systems [20], although only for spatially homogeneous conditions imposed by the DMFT assumption.
In this work, we develop a method that is capable of describing the spatio-temporal dynamics of condensed as well as non-condensed bosons in the case of strong Hubbard repulsion and, ideally, near the Mott-superfluid transition. We derive the equations of motion for the density matrix at a local site from the corresponding von Neumann equation, with second-order hopping processes included in the fluctuation correlations with the neighboring sites. Approximations leading to a time-local form of these correlators make it possible to identify these equations as time-dependent quantum master equations for the local density matrix in space. This approach allows us to study the expansion dynamics of interacting bosons in an optical lattice described by the BHM. Starting from a realistic initial spatial distribution of atoms with a Mott insulator at its core, surrounded by a superfluid ring, we observe the spatio-temporal dynamics to display a clear separation between a ballistically expanding condensate and a slow, diffusively expanding Mott insulator. These results are consistent with the experiment in Ref. [5], which cannot be reproduced by the standard Gutzwiller mean-field methods [15, 16]. The ballistic expansion velocity meets the Lieb-Robinson bound on information propagation [21], while the diffusive dynamics are corroborated by a slow expansion of the non-condensed atom cloud tproportional-$\propto\sqrt{t}$. Furthermore, we compute the von Neumann number entropy at local sites, which supports the expansion behavior and is relevant for the experiments with quantum gas microscopes [22].
The paper is organized as follows. In Section II we develop the method and derive the explicit form of the master equation for the density matrix of the BHM in Sections II.2 and II.3. The different expansion dynamics of condensate and non-condensed cloud are demonstrated in Section III, starting with the ground state of a lattice Bose gas in a harmonic trap, consisting of a Mott localized core surrounded by a condensed ring, as a realistic initial state of the expansion. Subsequently, we analyze in Sections IV and V the diffusive expansion of an initially purely Mott localized state and in Section V the ballistic expansion of an initially completely condensed gas. We conclude in Section VI with a brief summary of the merits and limitations, and possible future applications of the present method and of the specific results for the two-dimensional Bose-Hubbard model.
II Model and method
II.1 Density matrix approach to system dynamics coupled to a bath
The B-DMFT is not directly applicable to the expansion scenario, as the expanding cloud does not obey spatial translation symmetry. We do, however, adopt the DMFT concept of isolating a single site (the impurity) and embedding it in an effective bath that represents the remaining sites. The temporal dynamics of the isolated site will be treated accurately, assuming that the condensate and non-condensed densities of the surrounding sites are known. In the case of spatio-temporal expansion, however, these quantities will evolve in time and will have to be determined selfconsistently by treating each site as the dynamical impurity in the above sense. In this spirit, we first set up our method for a generic class of models with system-bath interaction, which can be described by the following Hamiltonian:
$\hat{H}=\hat{H}_{\rm S}+\hat{H}_{\rm B}+\hat{H}_{\rm SB}\,,$ (1)
where $\hat{H}_{\rm S,\,B}$ describes the system and the bath, respectively, and $\hat{H}_{\rm SB}$ the hybridization, which consists of the terms that act on both subsystems. The corresponding Liouville–von Neumann equation describing the dynamics of the system’s density matrix is given by
$\partial_{t}\hat{\rho}(t)=-i\left[\hat{H},\hat{\rho}(t)\right].$ (2)
A Gutzwiller-like mean-field approximation would allow for inter-site hopping only if there is a condensate amplitude on the neighboring sites and would, therefore, never delocalize an initial Mott state. To allow for transport of non-condensed particles by the hybridization Hamiltonian, we now switch to the interaction picture where the time evolution of an operator $\hat{O}$ from time $t_{0}$ to $t$ is governed by $\hat{O}^{I}(t)=\hat{U}^{\dagger}(t,t_{0})\hat{O}(t_{0})\hat{U}(t,t_{0})$ where $\hat{U}(t,t_{0})=\hat{U}_{\rm B}(t,t_{0})\hat{U}_{\rm S}(t,t_{0})$, with the unitary time-evolution operators in the system, $\hat{U}_{\rm S}(t,t_{0})$, and in the bath, $\hat{U}_{\rm B}(t,t_{0})$, defined as
$\hat{U}_{\rm S,\,B}(t,t_{0})=\hat{\mathcal{T}}e^{-i\int_{t_{0}}^{t}\mathrm{d}t^{\prime}\hat{H}_{\rm S,\,B}(t^{\prime})}\,,$ (3)
and the time-ordering operator $\hat{\mathcal{T}}$. Note that $\hat{H}_{\rm S}$ and $\hat{H}_{\rm B}$ of the isolated system and bath are a priori not time dependent, but will acquire an effective time dependence due to the dynamical partitioning into condensate and non-condensate particle densities, see below, which leads to the time-ordered integral in Eq. 3. The von Neumann equation for the density matrix $\hat{\rho}^{I}(t)$ in the interaction picture is given by
$\partial_{t}\hat{\rho}^{I}(t)=-i[\hat{H}_{\rm SB}^{I}(t),\hat{\rho}^{I}(t)]\,,$ (4)
which reads in integral form
[…]
By plugging Eq. 5 into Eq. 4, followed by tracing over the bath, we obtain the master equation for the system density-matrix to second order in the system-bath coupling,
$\partial_{t}\hat{\rho}^{I}_{\rm S}(t)=-\int_{t_{0}}^{t}\mathrm{d}t^{\prime}~{}{\rm Tr_{B}}\left[\hat{H}_{\rm SB}^{I}(t),\left[\hat{H}_{\rm SB}^{I}(t^{\prime}),\hat{\rho}^{I}(t^{\prime})\right]\right],$ (6)
where it is assumed that the system-bath coupling vanishes for $t\leq t_{0}$, so that ${\rm Tr_{B}}\left[\hat{H}_{\rm SB}^{I}(t),\hat{\rho}^{I}(t_{0})\right]=0$ [23]. Finally, by performing an inverse transformation of Eq. 6 from the interaction picture, the time-evolved system density matrix $\hat{\rho}_{\rm S}(t)$ in the Schrödinger picture is obtained, and the dynamics of the observables in the system can be studied.
Starting from Eq. 6, we will now apply this density matrix approach to study the spatio-temporal dynamics of the BHM as a generic example of interacting lattice-boson models.
II.2 Application to the Bose-Hubbard model
The BHM is described by the Hamiltonian [24, 25],
^$\hat{\mathcal{H}}_{\rm BHM}$ $=-J\sum_{\langle{\bf r},{\bf r^{\prime}}\rangle}\hat{b}_{\bf r^{\vphantom{\prime}}}^{\dagger}\hat{b}_{\bf r^{\prime}}^{\vphantom{\dagger}}+\sum_{\bf r}\hat{H}_{\bf r}^{\rm loc}$
${\rm with}~{}~{}\hat{H}_{\bf r}^{\rm loc}$ $=\left[\frac{U}{2}\hat{n}_{\bf r}(\hat{n}_{\bf r}-1)+V_{\bf r}\hat{n}_{\bf r}\right],$ (7)
where $\hat{b}_{\bf r}^{\vphantom{\dagger}}$, $\hat{b}_{\bf r}^{\dagger}$ and $\hat{n}_{\bf r}$ are the bosonic annihilation, creation, and number operators, respectively, at site ${\bf r}=(i_{x},i_{y})$ of a square lattice of size $L\times L$. $J$ is the nearest-neighbor hopping amplitude, $U$ the on-site interaction strength, and $V_{\bf r}$ an external potential, which can arise due to a magneto-optical trap or disorder, for example.
Figure 1: Schematic of the setup. (a) Sketch of the expansion of an interacting, bosonic cloud in an optical lattice. (b) and (c) show the zoomed-in spatial profiles of the particle density $N_{\bf r}(0)$ and of the superfluid density $|\Phi_{\bf r}(0)|^{2}$, respectively, for an initial Mott insulating core surrounded by a superfluid ring. The initial profile is computed using the Gutzwiller mean-field theory for $U/J=50$, $V_{0}/J=0.2$ and chemical potential $\mu/J=15$, see the text.
The Hamiltonian, Eq. 7, can be partitioned analogous to Eq. 1 as ^$\hat{\mathcal{H}}_{\rm BHM}=\hat{H}_{\rm S}+\hat{H}_{\rm B}+\hat{H}_{\rm SB}$, with
$\begin{split}\hat{H}_{\rm S}=\hat{H}_{\bf r}^{\rm loc},\quad\hat{H}_{\rm B}&=-J\sum_{\langle{\bf r^{\prime}},{\bf r^{\prime\prime}}\rangle\in({\bf r})}\hat{b}_{\bf r^{\prime}}^{\dagger}\hat{b}_{\bf r^{\prime\prime}}^{\vphantom{\dagger}}+\sum_{{\bf r^{\prime}}\neq{\bf r}}\hat{H}_{\bf r^{\prime}}^{\rm loc},\\ \hat{H}_{\rm SB}&=-J\sum_{{\bf r^{\prime}}\in{\bf R_{r}}}\hat{b}_{\bf r^{\prime}}^{\dagger}\hat{b}_{\bf r^{\vphantom{\prime}}}^{\vphantom{\dagger}}+{\rm H.c.}\,,\end{split}$ (8)
where an arbitrarily chosen site ${\bf r}$, which is fixed for the moment, represents the system and is coupled to the other lattice sites serving as a bath. Here, $({\bf r})$ denotes the set of all lattice sites except ${\bf r}$, and ${\bf R_{r}}$ the set of nearest neighbor sites of ${\bf r}$ [19]. By defining the condensate amplitude at site ${\bf r}$ as $\Phi_{\bf r}(t)=\langle\hat{b}_{\bf r}\rangle$ and the fluctuations as $\delta\hat{b}_{\bf r}(t)=\hat{b}_{\bf r}-\Phi_{\bf r}(t)$, the hopping term can be decomposed into mean-field (MF) parts in the system (S) and on the neighboring sites (B) and a fluctuation part (fl) as, $\hat{H}_{\rm SB}=\hat{H}_{\rm S}^{\rm MF}(t)+\hat{H}_{\rm B}^{\rm MF}(t)+\hat{H}_{\rm SB}^{\rm fl}(t)$, with
$\hat{H}_{\rm S}^{\rm MF}(t)$ $=-J\Phi_{\rm eff,{\bf r}}^{*}(t)\hat{b}_{\bf r}^{\vphantom{\dagger}}+{\rm H.c.}\,,$
$\hat{H}_{\rm B}^{\rm MF}(t)$ $=-J\sum_{{\bf r^{\prime}}\in{\bf R_{r}}}\Phi_{\bf r^{\vphantom{\prime}}}^{\vphantom{\dagger}}(t)\hat{b}_{\bf r^{\prime}}^{\dagger}+{\rm H.c.}\,,$ (9)
$\hat{H}_{\rm SB}^{\rm fl}(t)$ $=-J\sum_{{\bf r^{\prime}}\in{\bf R_{r}}}\delta\hat{b}_{\bf r^{\prime}}^{\dagger}(t)\delta\hat{b}_{\bf r^{\vphantom{\prime}}}^{\vphantom{\dagger}}(t)+{\rm H.c.}\,,$
where $\Phi_{\rm eff,\,{\bf r}}(t)=\sum_{\begin{subarray}{c}{\bf r^{\prime}}\in{\bf R_{r}}\end{subarray}}\Phi_{\bf r^{\prime}}(t)$ is the sum of the condensate amplitudes on the neighboring sites. Since $\hat{H}_{\rm S,\,B}^{\rm MF}(t)$ are local operators in the system or the bath, respectively, we can absorb these mean-field contributions in an effective system or bath Hamiltonian, respectively, and write the time-evolution operators of the uncoupled system and bath as (cf. Eq. (3))
$\hat{U}_{\rm S,\,B}(t,t_{0})=\hat{\mathcal{T}}e^{-i\int_{t_{0}}^{t}\mathrm{d}t^{\prime}\left(\hat{H}_{\rm S,\,B}+\hat{H}_{\rm S,\,B}^{\rm MF}(t^{\prime})\right)}$ (10)
Note that $\hat{H}_{\rm S}+\hat{H}_{\rm S}^{\rm MF}(t)$ corresponds to the Gutzwiller mean-field Hamiltonian [16], such that its dynamics are contained in the operator $\hat{U}_{\rm S}(t,t_{0})$. Following Eq. 6, the density matrix $\hat{\rho}_{\bf r}^{I}(t)$ of a local site ${\bf r}$ is time evolved with respect to $\hat{H}_{\rm SB}^{{\rm fl},\,I}(t)$ in the interaction picture, with its equation of motion given by
$\partial_{t}\hat{\rho}^{I}_{\bf r}(t)=-\int_{t_{0}}^{t}\mathrm{d}t^{\prime}~{}\mathrm{Tr}_{({\bf r})}\left[\hat{H}_{\rm SB}^{{\rm fl},\,I}(t),\left[\hat{H}_{\rm SB}^{{\rm fl},\,I}(t^{\prime}),\hat{\rho}^{I}(t^{\prime})\right]\right].$ (11)
We note in passing that the above partitioning is analogous to the cavity construction of DMFT, where one derives an effective action for the local (impurity) site after tracing over the bath sites, with the assumption that the self-energy is local [19]. In a similar spirit, spatial quantum correlations (beyond mean field) are neglected in our method, resulting in the factorization of the total density matrix as, $\hat{\rho}(t)=\prod_{\bf r}\hat{\rho}_{\bf r}(t)$. Dropping the fluctuation term $\hat{H}_{\rm SB}^{\rm fl}(t)$ in Eq. 9 would, thus, be equivalent to the single-site Gutzwiller mean-field theory [15, 16].
II.3 Derivation of the quantum master equation
Following Eq. 6, we now proceed to derive the time-evolution equation for the density matrix $\hat{\rho}_{\bf r}(t)$ at site ${\bf r}$ of the BHM. For convenience, we represent the field operators, $\hat{b}_{\bf r}^{\vphantom{\dagger}}$, $\hat{b}_{\bf r}^{\dagger}$, and the condensate amplitudes, $\Phi_{\bf r}^{\vphantom{*}}$, $\Phi_{\bf r}^{*}$, in Nambu space as
$\bm{b}_{\bf r}=\begin{pmatrix}\hat{b}_{\bf r}^{\vphantom{\dagger}}\\ \hat{b}_{\bf r}^{\dagger}\end{pmatrix},\quad\bm{\varphi}_{\bf r}=\begin{pmatrix}\Phi_{\bf r}^{\vphantom{*}}\\ \Phi_{\bf r}^{*}\end{pmatrix},$ (12)
respectively. The master equation for the density matrix (cf. Eq. 11) can then be written as
$\partial_{t}\hat{\rho}^{I}_{\bf r}(t)$ $=-J^{2}\int_{t_{0}}^{t}\mathrm{d}\bar{t}\sum_{{\bf r^{\prime}},{\bf r^{\prime\prime}}\in{\bf R_{r}}}\sum_{\alpha,\,\beta}\left(\left[b^{\dagger\,I}_{{\bf r^{\vphantom{\prime}}},\,\alpha}(t),b^{I{\vphantom{\dagger}}}_{{\bf r},\,\beta}(\bar{t})\hat{\rho}^{I}_{\bf r}(\bar{t})\right]\right.$
$-\left.\varphi_{{\bf r^{\vphantom{\prime}}}}^{\beta{\vphantom{\dagger}}}(\bar{t})\left[b^{\dagger\,I}_{{\bf r^{\vphantom{\prime}}},\,\alpha}(t),\hat{\rho}^{I}_{\bf r}(\bar{t})\right]\right)G^{(c)\alpha,\,\beta}_{{\bf r^{\prime},r^{\prime\prime}}}(t,\bar{t})+{\rm H.c.}\,,$ (13)
where $\alpha$ and $\beta$ denote the Nambu indices, and $G^{(c)\alpha,\,\beta}_{{\bf r^{\prime},r^{\prime\prime}}}(t,\bar{t})$ are the components of the connected part of the two-time correlator $\bm{G}_{{\bf r^{\prime}},\,{\bf r^{\prime\prime}}}^{(c)}(t,\bar{t})$ defined as
$G_{{\bf r^{\prime}},\,{\bf r^{\prime\prime}}}^{(c)\alpha,\,\beta}(t,\bar{t})=\mathrm{Tr}_{({\bf r})}\left\{{\delta}b^{I{\vphantom{\dagger}}}_{{\bf r^{\prime}},\,\alpha}(t)\delta b^{\dagger\,I}_{{\bf r^{\prime\prime}},\,\beta}(\bar{t})\hat{\rho}_{({\bf r})}^{I}(\bar{t})\right\}\,.$ (14)
We note that solving an integro-differential equation like Eq. 13 is challenging. To perform the retarded time integral, an exponential form of the bath correlation function has previously been assumed, see, for instance, Ref. [26]. In a similar spirit, we re-write Eq. 13 in Wigner relative and center-of-motion (CoM) time coordinates, $\tau=t-\bar{t}$ and $T=(t+\bar{t})/2$, respectively. For the present case (short-time correlations due to strong onsite repulsion $U$), we assume that in Wigner coordinates the two-time correlator is a sharply peaked function at time difference $\tau=0$ with an exponential decay in $\tau$,
$\bm{G}_{{\bf r^{\prime}},\,{\bf r^{\prime\prime}}}^{(c)}(t,\bar{t})\rightarrow\bm{G}_{{\bf r^{\prime}},\,{\bf r^{\prime\prime}}}^{(c)}(T,\tau)=\bm{G}_{{\bf r^{\prime}},\,{\bf r^{\prime\prime}}}^{(c)}(T,\tau=0)\,\mathrm{e}^{-\tau/\tau_{\rm corr}},$ (15)
with the correlation time $\tau_{\rm corr}\simeq 1/U$, while the rest of the integrand in Eq. 13 varies slowly as function of $\tau$, on a scale much longer than $\tau_{\rm corr}$. It can, thus, be pulled out of the $\tau$-integral and evaluated at equal times, $\tau=0$, so that the remaining part of the $\tau$-integral can be performed. Furthermore, since the density matrix factorizes in space, we obtain $\bm{G}_{{\bf r^{\prime}},\,{\bf r^{\prime\prime}}}^{(c)}(T,\tau=0):=\bm{G}_{{\bf r^{\prime}}}^{(c)}(t)\delta_{{\bf r^{\prime}},{\bf r^{\prime\prime}}}$. Transforming back to the Schrödinger picture re-introduces the on-site, coherent dynamics of the system by means of a von Neumann commutator (first term in Eq. 16), and we arrive at the master equation for $\hat{\rho}_{\bf r}(t)$ at the CoM time $t=T$ and $\tau=0$,
[…]
The last term in this equation reads,
$\mathcal{L}[\bm{b}_{\bf r}^{\vphantom{\dagger}}]$ $=\sum_{{\bf r^{\prime}}\in{\bf R_{r}}}\ \sum_{\alpha,\beta}{G}^{(c)\alpha,\beta}_{\bf r^{\prime}}(t)\left[2{b}_{{\bf r},\beta}^{\vphantom{*}}\hat{\rho}_{\bf r}(t){b}^{\dagger}_{{\bf r},\alpha}\right.$
$-\left.{b}^{\dagger}_{{\bf r},\alpha}{b}_{{\bf r},\beta}^{\vphantom{*}}\hat{\rho}_{\bf r}(t)-\hat{\rho}_{\bf r}(t){b}^{\dagger}_{{\bf r},\alpha}{b}_{{\bf r},\beta}^{\vphantom{*}}\right]\,,$ (17)
with the local, equal-time correlator,
${G}_{\bf r^{\prime}}^{(c)\alpha,\beta}(t)=$ $\mathrm{Tr}\left\{{\delta}{b}_{{\bf r^{\prime}},\alpha}^{\vphantom{\dagger}}(t)\delta{b}^{\dagger}_{{\bf r^{\prime}},\beta}(t)\hat{\rho}_{{\bf r}}(t)\right\}$ (18)
$=$ $\mathrm{Tr}\left\{{b}_{{\bf r^{\prime}},\alpha}^{\vphantom{\dagger}}{b}^{\dagger}_{{\bf r^{\prime}},\beta}\hat{\rho}_{{\bf r^{\prime}}}(t)\right\}-{\varphi}_{{\bf r^{\prime}},\alpha}^{\vphantom{*}}(t){\varphi}^{*}_{{\bf r^{\prime}},\beta}(t)\,,$
where the trace runs over the local Fock space at the nearest-neighbor site ${\bf r^{\prime}}$. The local system Hamiltonian which is renormalized by the Lamb-shift term (first term on the right-hand side of Eq. 16) is given by,
$\vphantom{\hat{H}}_{S}^{\rm MF}(t)=\hat{H}_{S}^{\rm MF}(t)+iJ^{2}\tau_{\rm corr}\sum_{{\bf r^{\prime}}\in{\bf R_{r}}}\left(\Phi_{\bf r^{\prime}}^{\vphantom{\dagger}}(t)\hat{b}_{\bf r}^{\dagger}-\Phi_{\bf r^{\prime}}^{*{\vphantom{\dagger}}}(t)\hat{b}_{\bf r}^{\vphantom{\dagger}}\right).$ (19)
The detailed derivation to arrive at this form of the Master equation is discussed in Appendix A. Once the density matrix $\hat{\rho}_{\rm r}(t)$ is obtained, one can compute any local, single-time expectation value as,
$\langle(\dots)\rangle={\rm Tr}\left[\hat{\rho}_{\bf r}(t)(\dots)\right]\,,$ (20)